OpenStudy (anonymous):

What is the eqution for a line passing through the points (0,3) and (3,5)? Answer in the form y=mx+c

OpenStudy (anonymous):

Read this - http://www.regentsprep.org/regents/math/algebra/ac1/eqlines.htm

OpenStudy (texaschic101):

use the slope formula and find the slope. slope(m) = (y2 - y1) / (x2 - x1) (0,3)....x1 = 0 and y1 = 3 (3,5)...x2 = 3 and y2 = 5 now we sub slope(m) = (5 - 3) / (3 - 0) can you finish and tell me the slope ?

OpenStudy (texaschic101):

when you want to participate, just let me know

OpenStudy (anonymous):

I thnk I got it now, thanks. m=2/3 and c is just the 3 from the given point (0,3)

OpenStudy (texaschic101):

lol...yes. But you already know it is correct. I commend you for your teaching.